Two particles each have a rest mass energy of 30 MeV and are traveling with a total of 10 MeV of kinetic energy. When they collide, they fuse into a new particle with a rest mass energy of 50 MeV. How much kinetic energy does the final particle possess?
Answer #1

Initial rest mass energy of two particles = 2*30 Mev = 60 Mev

Final rest mass energy of new particle = 50 Mev

Initial kinetic energy of two particles = 10 Mev

Let final kinetic energy of new particle = k Mev

Total energy before collision = total energy after collision.

60 + 10 = 50 +k

k = 20 Mev